# Relative Distance Guided Dynamic Partition Learning for Scale-Invariant UAV-View Geo-Localization

### v2

@@ IV-D Ablation Studies
$\alpha$=0.5 21.55 62.64 73.37 79.00 83.02 85.20 86.73 79.46 69.87 63.84 57.98 52.21 40.51 65.79
$\alpha$=1.5 23.69 63.57 74.08 80.36 83.98 85.44 86.31 80.11 72.49 67.29 62.14 56.59 46.20 67.86
$\alpha$=2.5 25.66 64.66 74.67 80.28 83.44 85.10 86.13 80.61 73.46 68.99 64.02 58.82 50.20 68.92
$\alpha$=3.3 24.84 65.39 75.61 80.70 83.41 85.01 85.72 81.87 75.34 71.19 67.17 63.42 55.64 70.40
$\alpha$=4.0 26.12 64.73 74.45 79.74 82.63 84.27 84.87 80.67 74.73 71.35 67.83 64.42 56.77 70.19
Number of partitions. The number of parts $N_{SPS}$ is a key hyper-parameter in our DGDPL. As shown in Table VI, we report the quantitative results of DGDPL with different $N_{SPS}$ on the three test-sets. As the number of parts increases, the retrieval accuracy of DGDPL exhibits a gradual enhancement, particularly in scenarios involving inconsistent scales. For example, when $N_{SPS}$ increases from 1 to 3, DGDPL improves R@1 by 12.58%, 18.87% and 30.96% in the standard test set, Short($\Delta P$=-100) and Long($\Delta P$=-100), respectively. Further increasing $N_{SPS}$, the retrieval performance of DGDPL tends to be stable with a slight improvement, because 3 or 4 partitions are enough to exploit rich contextual information. As the basic component of DGDPL, our SPS divides high-level features into fixed-size fine-grained partitions. To thoroughly demonstrate the anti-scaling robustness of SPS, we also report results for LPN [24] and SPS under various number of partitions, as shown in Fig. 9. It is evident that our SPS exhibits comparable performance to the LPN on the standard test set, while outperforming LPN significantly in scale-inconsistent scenarios. When SPS meets DGAS and SGRS, model (i.e., DGDPL) with various $N_{SPS}$ achieves significant performance improvement in each test set.
TABLE VI: performance comparison of DGDPL with different number of parts. $N_{SPS}$ denotes the number of parts generated by $SPS$.
@@
4 84.19 86.36 65.39 69.41 63.42 68.05
Figure 9: Comparison of LPN, SPS and DGDPL with different numbers of parts. The experiments were conducted on the standard test set, Short($\Delta P$=-100) and Long($\Delta P$=100).
input image size. A small training size compresses the fine-grained information of the input image, thus compromising the discriminative representation learning. We selected five input image resolutions to investigate the impact of input image size on model performance. To guarantee that the partition resolution remains positive, we adjust the scaling factor $alpha$ in accordance with variations in the input image resolution. As shown in Table VII, when the input image resolution changes from 224 to 512, model performance improves continuously on multiple datasets. Furthermore, with the image resolution increasing from 320 to 512, the gain of DGDPL is 3.11% on the standard test set, while the improvement is 8.88% and 15.74% on Short($\Delta P$=-100) and Long($\Delta P$=100), respectively. We also present the results of LPN [24], the most relevant study, under different input picture sizes, as depicted in Fig. 10. With the increase of image resolution, the accuracy of LPN and DGDPL is equivalent in the standard test set, but the performance gap between them gradually widens in extreme scenes. It can also be found that the performance of LPN is unstable in Short($\Delta P$=-100) and Long($\Delta P$=100). This indicates that DGDPL can extract the rich fine-grained information contained in high-resolution images.
TABLE VII: performance comparison of DGDPL with different input image size.
@@
V Conclusions
In this paper, we study the UAV-view geo-localization in scale-inconsistent scenarios, and propose a simple and effective part-based representation learning framework called DGDPL. Specifically, we introduce a square partition strategy to capture fine-grained features while retaining the global structure. Adhering to the rule that objects in UAV-views appear larger when closer and smaller when farther, we further introduce a distance-guided adjustment strategy. Based on the distance ratio of cross-views, we dynamically adjust the partition size of the UAV branch to promote the content consistency of partition pairs, thereby achieving accurate cross-view matching against scale variations. Moreover, we explore a saliency-guided refinement strategy to refine part-level features into salient features and background features, further improving cross-view image matching performance without compromising the robustness of our model to scale variations. Extensive experiments on the University-1652 demonstrate that the robustness of our method in various scale-inconsistent scenarios. Ablation experiments show that DGDPL reveals satisfactory compatibility with various backbone networks. In the future, we plan to collect a real-world dataset containing UAV pose information to drive real-world cross-view matching, such as UAV-view geo-localization and UAV self-positioning.

### v3

@@ IV-D Ablation Studies and Further Discussion
$\alpha$=2.0 21.55 62.64 73.37 79.00 83.02 85.20 86.73 79.46 69.87 63.84 57.98 52.21 40.51 65.79
$\alpha$=4.0 21.71 61.99 73.60 79.78 83.22 85.29 86.83 79.93 71.70 65.67 59.70 53.61 42.04 66.54
$\alpha$=6.0 23.69 63.57 74.08 80.36 83.98 85.44 86.31 80.11 72.49 67.29 62.14 56.59 46.20 67.86
$\alpha$=8.0 23.24 63.36 73.84 80.05 83.37 85.36 86.54 80.35 72.71 68.38 63.38 58.02 47.59 68.17
$\alpha$=10.0 25.66 64.66 74.67 80.28 83.44 85.10 86.13 80.61 73.46 68.99 64.02 58.82 50.20 68.92
$\alpha$=12.0 23.52 63.40 73.88 79.68 82.98 85.28 86.31 81.18 74.85 71.18 67.08 62.79 54.68 69.75
$\alpha$=14.0 25.68 64.89 74.95 80.20 83.54 85.09 86.16 81.88 75.95 72.24 68.47 64.08 55.48 70.66
$\alpha$=16.0 26.12 64.73 74.45 79.74 82.63 84.27 84.87 80.67 74.73 71.35 67.83 64.42 56.77 70.19
Number of partitions. The number of parts $N$ is a key hyper-parameter in SaLPN. As shown in Table VII, we report the quantitative results of SaLPN with different $N$ on the three test-sets. As the number of parts increases, the retrieval accuracy of SaLPN exhibits a gradual enhancement, particularly in scenarios involving inconsistent scales. For example, when $N$ increases from 1 to 3, SaLPN improves R@1 by 12.58%, 18.57% and 30.96% in the standard test set, Short ($\Delta P$=-100) and Long ($\Delta P$=-100), respectively. Further increasing $N$, the retrieval performance of SaLPN tends to be stable with a slight improvement, because 3 or 4 partitions are enough to exploit rich contextual information. As the basic component of SaLPN, SPS divides high-level features into fixed-size fine-grained partitions. To thoroughly validates the anti-scaling robustness of SPS, we also report results for LPN [25] and SPS under various number of partitions, as shown in Fig. 9. It is evident that SPS exhibits comparable performance to LPN on the standard test set, while outperforming LPN significantly in scale-inconsistent scenarios. When SPS is combined with HAAS and SGRS, model (i.e., SaLPN) with various $N$ achieves significant performance improvement in each test set.
TABLE VII: performance comparison of SaLPN with different number of parts. $N$ denotes the number of parts generated by $SPS$.
@@
4 84.37 86.51 64.89 68.86 64.08 68.75
Figure 9: Comparison of LPN, SPS and SaLPN with different numbers of parts. The experiments were conducted on the standard test set, Short ($\Delta P$=-100) and Long ($\Delta P$=100).
Input image size. A small training size compresses the fine-grained information of the input image, thus compromising the discriminative representation learning. We select five input image resolutions to investigate the impact of input image size on model performance. To guarantee that the partition resolution remains positive, we adjust the partition adjustment factor $\alpha$ in accordance with variations in the input image resolution. As shown in Table VIII, when the input image resolution changes from 224 to 512, model performance improves continuously on multiple datasets. Furthermore, with the image resolution increasing from 320 to 512, the gain of SaLPN is 3.29% on the standard test set, while the improvement is 8.38% and 16.40% on Short ($\Delta P$=-100) and Long ($\Delta P$=100), respectively. We also present the results of LPN [25], the most relevant study, under different input picture sizes, as depicted in Fig. 10. With the increase of image resolution, the accuracy of LPN and SaLPN is equivalent in the standard test set, but the performance gap between them gradually widens in extreme scenes. It can also be found that the performance of LPN is unstable in Short($\Delta P$=-100) and Long($\Delta P$=100). This indicates that SaLPN can extract the rich fine-grained information contained in high-resolution images.
TABLE VIII: performance comparison of SaLPN with different input image size.
@@
V Conclusion
In this paper, we study the UAV-view geo-localization in scale-inconsistent scenarios, and propose a simple and effective part-based representation learning framework called SaLPN. Specifically, we introduce a square partition strategy to capture fine-grained features while retaining the global structure. Adhering to the rule that objects in drone-views appear larger when closer and smaller when farther, we further introduce a height-aware adjustment strategy. Based on the height ratio of cross-view images, we dynamically adjust the partition size of the drone branch to promote the content consistency of partition pairs, thereby achieving accurate cross-view matching against scale variations. Moreover, we explore a saliency-guided refinement strategy to refine part-level features into salient features and background features, further improving cross-view image matching performance without compromising the robustness of our model to scale variations. Extensive experiments on University-1652 and SUES-200 validate that the effectiveness and robustness of our method in various scale-inconsistent scenarios. Ablation experiments show that SaLPN reveals satisfactory compatibility with various backbone networks. In the future, we plan to collect a real-world dataset containing drone pose information to drive real-world cross-view matching, such as drone-view geo-localization and drone self-positioning.